Mustansir is the owner of small pottery firm. He has hired one helper at Rs. 15,500/= per month,

pays annual rent of 144,000/=, spends around 19,000/= on materials. He has 400,000/= of his own

funds invested in equipment like pottery wheels, kilns and so forth that could lead him to earn

50,000 per year if alternatively, invested. He was also offered by his competitor to work as a potter


and earn 60,000/= per month. He estimated his entrepreneurial skills worth Rs. 53,000 per year.

Although total annual revenue from the pottery sales is 9,650,000. Calculate the accounting and

economic profit of Mustansir’s pottery firm and comment on the situation that whether he should

continue doing the business or shut down.


1
Expert's answer
2021-06-15T10:14:58-0400

Solution


Accounting profit=

Total revenue - explicit cost

Explicit cost=hired helper

15500"\times12=186000"

Annual rent= 144000

Materials =19000

Therefore explicit costs amounts to

"186000+144000+19000 =" "349000"

Accounting profit="9650000 - 349000= 9,301,000"


Economic profit= Total revenue -(explicit + implicit costs)


Implicit cost= foregone salary(60000"\times12)= 720000"

Normal profit=53000

Foregone interest=50000


Implicit cost therefore amounts to

"720000+53000+50000 =" "823000"

Economic "=9301000-823000\n\n=8,478,000"

The business receives more profits compared to it's input therefore Mustansir should continue doing the business.